OpenStudy (sugardemonhunters):

SIMPLE INTEREST- please help! :) How much simple interest is earned on $2000 at 4% interest for 6 months? Give your answer in dollars and cents. Round your answer to the nearest hundredth. Simple Interest I = p • r • t

OpenStudy (jadeishere):

Are you doing Exponential Growth and Decay functions, or Compound Interest right now in your math class?

OpenStudy (sugardemonhunters):

Compound :)

OpenStudy (jadeishere):

Okay :D Are you required to use the "formula" in your question, if so you just plug in the values p = starting amount r = rate of growth (percentage) t = time period

OpenStudy (sugardemonhunters):

So do I multiply all of them?

OpenStudy (jadeishere):

Let me work it out on paper then I'll show you >.<

OpenStudy (sugardemonhunters):

Thanks :D

OpenStudy (jadeishere):

Okay, yeah, sorry xD I was overthinking this for a moment. The starting amount is at $2000, the simple interest rate is at 4% and the time is at 6 months. Use the formula, l = p * r * t l = 2000 * 0.04 * 6

OpenStudy (jadeishere):

So yeah, multiply them together :)

OpenStudy (sugardemonhunters):

would the answer be the final answer?

OpenStudy (jadeishere):

I didn't multiply them together, they still need to be multiplied, so in a calculator plug in the values, and then that should be your answer. It says to round to the nearest tenth, so if the value is, let's say.. 460.89, then the answer would be $460.90

OpenStudy (sugardemonhunters):

Oh yeah I know but like in future questions like this all i'd have to do is multiply them? Thank you so much by the way~

OpenStudy (jadeishere):

Your welcome, and to answer your question, No, not necessarily. When it comes to Simple Interest, yes, this would be the usual way to solve them, but when coming across exponential growth, decay and Compound Interest, there are specific formulas for the equations.

OpenStudy (sugardemonhunters):

okay! thanks again

OpenStudy (jadeishere):

No problem :D